OpenStudy (anonymous):

Rewrite cube root of (27x-81) - 5 to make it easy to graph using a translation. Describe the graph. How do I rewrite it?

OpenStudy (anonymous):

@mathmale can you help me out real quick?

OpenStudy (ericgai):

\[\sqrt[3]{27x-81}-5=\sqrt[3]{27}\sqrt[3]{x-3}-5=3\sqrt[3]{x-3}-5\]